An object is moving on a straight line with a constant acceleration. At t=0 its velocity is 20 m/s.it stops after 5 s. If the mass is 3 kg then calculate the work done by net force
1
Expert's answer
2020-09-23T09:00:20-0400

Use the work-energy theorem:


"W=\frac{mv^2_f}{2}-\frac{mv^2_i}{2}"

Since the final velocity s 0:


"W=-\frac{mv^2}{2}=-\frac{3\cdot20^2}{2}=-600\text{ J}."